The line joining the point 𝐴 (4,6) π‘‘π‘œ 𝐡 (𝑏,βˆ’3) has a gradient of 3. Find the value of 𝑏.


1
Expert's answer
2022-02-06T16:51:53-0500
"m=gradient=\dfrac{y_B-y_A}{x_B-x_A}"


Given "A(4, 6), B(b, -3), m=gradient=3"


Substitute


"3=\dfrac{-3-6}{b-4}"

"b-4=-3"

"b=1"